OpenStudy (alystar):

Can someone give me a refresher on this Suppose 𝑀 and π‘₯ vary inversely and π‘₯ = 1.25 when 𝑀 = 8. Find π‘₯ when 𝑀 = 15.

OpenStudy (zzr0ck3r):

\(w=\dfrac{k}{x}\) \(8=\dfrac{k}{1.25}\) can you solve for \(k\)?

OpenStudy (alystar):

10

OpenStudy (zzr0ck3r):

so \(w=\dfrac{10}{x}\) Now plug in \(15\) for \(w\) and solve for \(x\)

OpenStudy (alystar):

2/3

OpenStudy (alystar):

Thanks, so much!

OpenStudy (zzr0ck3r):

np